OpenStudy (anonymous):

whats the expression for 1/a + 1/b ? and the expression 1/a+b ?

OpenStudy (anonymous):

\[\frac{a\cdot b}{a+b}\]

OpenStudy (anonymous):

You can leave \[\frac{1}{a+b}\] like it is.